OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form.

OpenStudy (lalaly):

parallel lines have the same slope the slope here y=mx+b m=-1 and it passes through (-4,2) so substitute in the equation y=mx+b to find b 2=-1(-4)+b b=2-4=-2 so the equation is y=-x-2

OpenStudy (lalaly):

and lol at what u deleted

OpenStudy (lalaly):

xD

OpenStudy (anonymous):

Nono... lalaly. He wants it in intercept form. So it should be:- Since the line is parallel to the line given. The gradient of the line is -1. Hence, \[(y-2)\div(x+4) = -1\]\[y-2 = -x-4\]\[x+y = -2\]\[-x/2 - y/2 = 1\]

OpenStudy (nilankshi):

hmmmmmmm